# Feshbach-Villars Formalism for a Spin-1/2 Particle in Curved Spacetime

### v1

@@ IV Conclusion
$\gamma^{\mu}\gamma^{\nu}=g^{\mu\nu}+\sigma^{\mu\nu},\quad\text{where }\sigma^{\mu\nu}=\frac{1}{2}\left[\gamma^{\mu},\gamma^{\nu}\right].$ (51)
@@
$-\gamma^{\mu}\gamma^{\nu}\mathcal{D}_{\mu}\mathcal{D}_{\nu}=-g^{\mu\nu}\mathcal{D}_{\mu}\mathcal{D}_{\nu}-\sigma^{\mu\nu}\mathcal{D}_{\mu}\mathcal{D}_{\nu}.$ (52)
@@
$[\mathcal{D}_{\mu},\mathcal{D}_{\nu}]=\frac{1}{4}R_{\mu\nu\lambda\sigma}\gamma^{\lambda\sigma}-iqF_{\mu\nu}.$ (53)
@@
Adding the mass term $-m^{2}$:
@@
Here, the derivation is exact under the assumptions of: Minimal coupling to gravity and electromagnetism, Metric compatibility $(\nabla_{\mu}g_{\alpha\beta}=0)$ and Torsion-free spacetime. It is important to mention that (58) was originally derived by Schrodinger (17, 2025).

### v2

@@ VII Data Availability Statement
[…]
@@
$\gamma^{\mu}\gamma^{\nu}\mathcal{D}_{\mu}\mathcal{D}_{\nu}=-g^{\mu\nu}\mathcal{D}_{\mu}\mathcal{D}_{\nu}-\sigma^{\mu\nu}\mathcal{D}_{\mu}\mathcal{D}_{\nu}.$ (131)
@@
$\mathcal{D}_{\mu},\mathcal{D}_{\nu}]=\frac{1}{4}R_{\mu\nu\lambda\sigma}\gamma^{\lambda\sigma}-iqF_{\mu\nu}.$ (132)
@@
Adding the mass term $m^{2}$:
@@
Here, the derivation is exact under the assumptions of: Minimal coupling to gravity and electromagnetism, Metric compatibility $(\nabla_{\mu}g_{\alpha\beta}=0)$ and Torsion-free spacetime. It is important to mention that (137) was originally derived by Schrodinger Schrodinger1932.
Appendix B Proof of Eq. (89)
The FV formalism rewrites the wave function as:
$\psi=\phi+\chi$ (138)
and relates the time derivative to the difference:
$i\partial_{t}\psi=m(\phi-\chi)$ (139)
This allows you to express $\phi$ and $\chi$ in terms of $\psi$ and its time derivative:
$\phi=\frac{1}{2}\left(\psi+\frac{i}{m}\partial_{t}\psi\right)$ (140)
$\chi=\frac{1}{2}\left(\psi-\frac{i}{m}\partial_{t}\psi\right)$ (141)
Given:
$i\partial_{t}\phi=\frac{m}{2}\phi-\frac{1}{2m}\mathcal{D}\chi$ (142)
$i\partial_{t}\chi=-\frac{m}{2}\chi-\frac{1}{2m}\mathcal{D}\phi$ (143)
where
$\mathcal{D}=\partial_{r}^{2}+\frac{1}{r}\partial_{r}+\frac{1}{\alpha^{2}r^{2}}\partial_{\varphi}^{2}$ (144)
Differentiate the first equation with respect to time and use the second to eliminate $\chi$:
  • Start with the first equation:
    $i\partial_{t}\phi=\frac{m}{2}\phi-\frac{1}{2m}\mathcal{D}\chi$ (145)
  • Take the time derivative:
    $i\partial_{t}(i\partial_{t}\phi)=i\partial_{t}\left(\frac{m}{2}\phi-\frac{1}{2m}\mathcal{D}\chi\right)$ (146)
    $\partial_{t}^{2}\phi=\frac{m}{2}i\partial_{t}\phi-\frac{1}{2m}\mathcal{D}i\partial_{t}\chi$ (147)
  • Substitute $i\partial_{t}\chi$ from the second equation:
    $i\partial_{t}\chi=-\frac{m}{2}\chi-\frac{1}{2m}\mathcal{D}\phi$ (148)
  • Plug into the previous expression:
    $\partial_{t}^{2}\phi=\frac{m}{2}i\partial_{t}\phi-\frac{1}{2m}\mathcal{D}\left(-\frac{m}{2}\chi-\frac{1}{2m}\mathcal{D}\phi\right)$ (149)
    $\partial_{t}^{2}\phi=\frac{m}{2}i\partial_{t}\phi+\frac{m}{2}\frac{1}{2m}\mathcal{D}\chi+\frac{1}{4m^{2}}\mathcal{D}^{2}\phi$ (150)
    $\partial_{t}^{2}\phi=\frac{m}{2}i\partial_{t}\phi+\frac{1}{2}\mathcal{D}\chi+\frac{1}{4m^{2}}\mathcal{D}^{2}\phi$ (151)
Now, recall $\psi=\phi+\chi$. Using the above and similar steps for $\chi$, you can combine to get a second-order equation for $\psi$.
Following the FV formalism for spin-1/2 in curved spacetime, the second-order equation for $\psi$ in your geometry (static cosmic string, 2D) is:
$\left[\partial_{t}^{2}-\mathcal{D}+m^{2}\right]\psi=0$ (152)
where
$\mathcal{D}=\partial_{r}^{2}+\frac{1}{r}\partial_{r}+\frac{1}{\alpha^{2}r^{2}}\partial_{\varphi}^{2}$ (153)
This matches the generalized Klein-Gordon equation obtained by squaring the Dirac operator and applying the FV transformation.